OpenStudy (anonymous):

one more help please. some difficult DEs which I can't figure out. dx + (y^2x)dy = (x^2 y) dy

OpenStudy (anonymous):

it is : dx + (y^2 *x)dy = (x^2 *y) dy

OpenStudy (amistre64):

i wonder: \[\frac{dx}{dy}+xy^2=x^2y\] \[x^{-2}x'+y^2x^{-1}=y\] \[z=x^{-1}:~z^{-1}=x;~z^{2}=x^{-2};~-z^{-2}z'=x';~\] \[z^2(-z^{-2}z')+y^2z=y\] \[-z'+y^2z=y\] \[z'-y^2z=-y\] justwondering, i got no idea if its right tho

OpenStudy (amistre64):

\[e^{-\frac{1}{3}y^3}z'-y^2e^{-\frac{1}{3}y^3}z=-ye^{-\frac{1}{3}y^3}\] \[e^{-\frac{1}{3}y^3}z=-\int ye^{-\frac{1}{3}y^3}dy\] \[z=-~e^{\frac{1}{3}y^3}\int ye^{-\frac{1}{3}y^3}dy\]

OpenStudy (amistre64):

i doubt that was right; unless we were spose to go into the Gamma function :/

OpenStudy (amistre64):

any ideas on how to start lilMiss ??

OpenStudy (anonymous):

sorry. poor internet connection here. well, what i was thinking was to get the integrating factor of the equation.

OpenStudy (anonymous):

just the same as what you did, i guess?

OpenStudy (amistre64):

the twp dys have me iffy; i have never had to deal with them in that state

OpenStudy (amistre64):

dx + (y^2x)dy = (x^2 y) dy dx = (x^2y - y^2x)dy dx/dy = x^2y - y^2x

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=dx%2Fdy+%3D+x%5E2y+-+y%5E2x i run into a gamma function at every turn :/

OpenStudy (anonymous):

help! integrate this please. : int. e^(-1/3 *y^3)dy

OpenStudy (anonymous):

i think there is no closed form for that integral ---------------------------------------------------------------------- \[M\text{d}x+N\text{d}y=0\]another method is multiplynig by \(x^a.y^b\) and searching for \(a\) and \(b\) for which\[\frac{\partial (x^a.y^b.M)}{\partial y}=\frac{\partial (x^a.y^b.N)}{\partial x}\]which turnes the equation to an exact diff equation let me try it

OpenStudy (anonymous):

no its not working im wondering where these equations come from...this one and prev are very curious...

OpenStudy (anonymous):

i already got this one, using bernoulli's equation with respect to x, are you familiar wit that? my only problem is how am i going to integrate e^(-1/3 *y^3)dy

OpenStudy (anonymous):

\[\large \int e^{-\frac{y^3}{3}} dy\]?

OpenStudy (anonymous):

but there is no closed form for this integral...

OpenStudy (anonymous):

why? i'll try it in wolfram alpha

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

will i pass my midterm? :(

OpenStudy (anonymous):

sure u will...dont think of these 2 prev problems

OpenStudy (anonymous):

they were not normal

OpenStudy (anonymous):

aw. thank you so much. really really thank you for everything that you've done just so i could have answers. one medal isn't enough, i guess. :)

OpenStudy (anonymous):

welcome...:)